# Strain effects on the electronic properties of a graphene wormhole

### v1

@@ Abstract
In this work, we explore the strain and curvature effects on the electronic properties of a curved graphene structure, called the graphene wormhole. The electron dynamics is described by a massless Dirac fermion containing position–dependent Fermi velocity. In addition, the strain produces a pseudo–magnetic vector potential to the geometric coupling. For an isotropic strain tensor, the decoupled components of the spinor field exhibit a supersymmetric (SUSY) potential, depending on the centrifugal term and the external magnetic field only. In the absence of a external magnetic field, the strain yields to an exponential damped amplitude, whereas the curvature leads to a power–law damping of the wave function. The spin–curvature coupling breaks the chiral symmetry between the upper and the lower spinor component, which leads to the increasing of the wave function on either upper or lower region of the wormhole, i.e., depending on the spin number. By adding an uniform magnetic field, the effective potential exhibits an asymptotic quadratic profile and a spin–curvature barrier near the throat. As a result, the bound states (Landau levels) are confined around the wormhole throat showing an asymmetric and spin–dependent profile.
@@ I Introduction
Two dimensional materials, such as graphene geim , silicene silicene and phosphorene phosphorene , have been the subject of intense investigations due to their outstanding properties. Beyond the remarkable mechanical katsnelson and electronic properties Novoselov2004 ; electronic , graphene can also be seen as a table–top laboratory for relativistic physics. Indeed, since the conduction electrons are effectively described as massless Dirac fermions, relativistic effects such as zitterbewegung zitter , Klein tunneling klein and atomic collapse collapse have been observed. Since the graphene layer can assume a curved shape, the curvature effects might lead to new interesting relativistic effects, such as the Hawking–Unruh effect hawking ; hawking2 .
@@
The study of a Dirac fermion confined into a two dimensional surface was initially addressed in Ref.BJ and further developments were provide afterwards diracsurface ; diracsurface2 ; diracsphere . For a relativistic fermion intrinsically living on a curved surface, a physical realization was found for conducting electrons on two dimensional carbon–based structures, as the fullerenes diracintrinsic ; diracintrinsic1 , carbon nanotubes saito and graphitic cones cone . In graphene, the massless Dirac equation in curved spaces was studied in a variety of shapes, such as the localized gaussian bump contijo , the cone furtado , a helical graphene ribbon atanasov ; watanabe , a corrugated plane corrugated , a Möbius ring mobius1 ; mobius2 , a torus ozlem ; Yesiltas:2021crm among others. The surface curvature produces a spin–curvature coupling which leads to a geometric Aharonov–Bohm–like effect geometricphase , a modified spin–orbit coupling geometricsoc ; geometricsoc2 and a geometric spin–Hall effect geometricmonopole .
@@
Besides the curvature, the deformations of the graphene layer modify the effective Dirac fermion dynamics as well, producing the so–called pseudo–magnetic fields ribbons . This vector potential steams from the strain tensor defined by the deformations of the graphene layer and the pseudo–magnetic term comes from the coupling to the Dirac fermion; it is similar to the minimal coupling to a magnetic field gaugestrain . The strain applied to graphene can mimic a strong magnetic field strainstrong and lead to important applications strainappli . From the strain tensor, an effective Hamiltonian for the Dirac fermion was derived using the tight–binding approach vozmediano , containing an anisotropic and position–dependent Fermi velocity and a pseudo–magnetic strain vector in the continuum limit. Since the strain may contain both in–plane and out–of–plane components, the effective Dirac Hamiltonian was extended in order to encompass all the stretching and bending effects vozmediano2 . In addition, the quantum field interaction of the effective Dirac fermion and the strain was discussed in sinner ; gaugegrapheneqft .
@@
An interesting curved graphene structure is the so–called graphene wormhole, where two flat graphene layers are connected by a carbon nanotube wormhole . Since its shape (cylinder) has a non–vanishing mean curvature, the discontinuity of the curvature at the graphene–nanotube junction leads to modifications of the energy spectrum and the possibility of localized states near to it graphenejunction . Although, the Dirac fermions on the upper and lower layers are free states (non–normalizable), the curvature of the nanotube allows the existence of normalizable zero–modes confined at the radius of the wormhole picak ; wormhole3 . In order to avoid the discontinuity at the junction, a smooth graphene wormhole was proposed considering a continuum and asymptotic flat catenoid surface dandoloff ; euclides ; deSouza:2022ioq . The negative curvature of the catenoid leads to a repulsive spin–curvature coupling near the wormhole throat, allowing only the zero–mode as a localized state around the throat wormhole4 ; ozlem2 ; wormhole5 .
@@
In this work, we consider the effects of the curvature and the strain on the effective Dirac fermion living in a catenoid–shaped graphene wormhole. We extend the effective Hamiltonian obtained in vozmediano to a curved surface, introducing the usual spin–connection coupling. We explore the different effects driven by the curvature, isotropic strain and an external magnetic field. Since the surface is asymptotic flat, the lattice deformation which yields the curvature and strain should be concentrated around the throat. The strain leads to a vector potential along the surface meridian, whereas the spin–curvature coupling points in the angular direction. Moreover, the strain vector potential provides an exponential damping of the wave function, whereas the curvature leads to a power–law decay. By adopting the so-called supersymmetric quantum mechanic-like approach ozlem2 , the spinor components exhibit a chiral symmetry breaking. Indeed, the upper component has its probability density enhanced near the wormhole throat in the upper layer, whereas the lower component is enhanced in the lower layer. The ground state zero mode also exhibits this chiral behaviour, since it is exponentially damped either in the upper or in the lower layer depending on the total angular momentum. By applying an uniform magnetic field, the Landau levels are also modified by the curved geometry and strain, leading to asymmetric localized states near the throat.
@@
This work is organized as the follows. In the section (II), we provide a brief review on the catenoid-shape graphene wormhole geometry. In the section (III) we present the effective Hamiltonian containing the strain, curvature and external magnetic field interactions. The section (IV) is devoted to the symmetries of the effective Hamiltonian and in the section (V) we employ the SUSY-QM approach in order to investigate the effects of each interaction. Finally, additional discussion and perspectives are outlined in the section (VI).
@@ II Graphene wormhole geometry
In this section, we define the graphene wormhole surface and describe some of its most important properties. We consider a smooth surface connecting an upper to the lower layer (flat planes). For this purpose, we choose a catenoid shaped surface. In other words, the catenoid surface can be describe in coordinates by euclides ; ozlem2
@@
Remember that they modify the Fermi velocity by turning it into a position dependet configuration. Moreover, from the dreinbeins, we can defined the moving frame $\theta^{a}=e^{a}_{\mu}\mathrm{d}x^{\mu}$, where, in the catenoid, take the form $\theta^{0}=\mathrm{d}t$, $\theta^{1}=\mathrm{d}u$, $\theta^{2}=\sqrt{R^{2}+u^{2}}\mathrm{d}\phi$. From the torsion–free condition, i.e., $T^{a}=\mathrm{d}\theta^{a}+\omega^{a}_{b}\wedge\theta^{b}=0$, the only non–vanishing one–form connection coefficient $\omega^{a}_{b}=\Gamma_{cb}^{a}\theta^{c}$ is given by
@@
The curvature 2–form $R^{a}_{a}=d\omega^{a}_{b}+\omega^{a}_{c}\wedge\omega^{c}_{b}$ has only non–vanishing component, namely $R^{2}_{1}=-\frac{R^{2}}{(R^{2}+u^{2})^{2}}\theta^{2}\wedge\theta^{1}$. Accordingly, the gaussian curvature $K=\delta^{ab}R_{ab}$ has the form
@@ III Strain Hamiltonian
After a brief review of the main geometric properties of the graphene wormhole, let us now discuss the effective Hamiltonian containing the strain and curvature effects on the electron. We follow closely the Ref.(vozmediano ) where the most general effective Hamiltonian was derived. The effective Hamiltonian for a Dirac fermion constrained to a flat surface under influence of the strain and an external magnetic field was found in Ref. (vozmediano ). We propose a generalization of the effective Hamiltonian of Ref.(vozmediano ) in the continuum limit for curved surfaces in the form
@@
The definition of the curved sigma matrices employed in Eq.(12) ensures that these matrices do not depend on the particular choice of coordinates of the surface (surface covariance). It is worthwhile to mention that the definitions of the velocity tensor in Eq.(24 and the curved Pauli matrices in Eq.(12) lead to a position and direction dependent Dirac kinetic term $H_{1}=v_{i}^{j}\sigma^{i}\partial_{j}$.
@@
The curved $\gamma^{\mu}$ matrices are related to the flat $\gamma^{a}$ ones by the dreinbeins $e^{a}_{\mu}$, i.e., $\gamma^{\mu}=e^{\mu}_{a}\gamma^{a}$. The dreinbeins are defined as $g_{\mu\nu}=e_{\mu}^{a}e_{\nu}^{b}$. In $(2+1)$ dimension, we can adopt the following representation to the flat Dirac $\gamma^{a}$ matrices $\gamma_{0}=\sigma_{3}$, $\gamma_{1}=-i\sigma_{2}$ and $\gamma_{2}=-i\sigma_{1}$ mobius2 ; ozlem ; ozlem2 . Thus, the curved Dirac matrices on the wormhole graphene surface have the form
@@
From the connection 1–form in Eq.(7), only $\omega^{2}_{1}$ is non zero. Thus, the only non–vanishing component of the spinor connection $\Omega_{\mu}$ is
@@
Note that, since $-\infty<u<\infty$, the geometric spinor potential in Eq.(18) is an odd function under parity. This parity violation does not occur for the Dirac fermion in a flat plane diracplanar or the graphitic cone furtado . Moreover, for $R=0$ or for $R\neq 0$ and $u\rightarrow\pm\infty$, the geometric connection is constant, as found for conic surfaces furtado . It is worth mentioning that, due to the resemblance of the spinor and gauge field coupling, the spinor potential is sometimes interpreted as a kind of a pseudo–magnetic potential steaming from the curved geometry contijo .
Furthermore, the strain produces two different potentials on the Dirac electron. The first potential, steaming from the strain in Eq.(11), is a vector potential, whereas the second one in Eq.(18) is a spinorial potential depending on the surface connection. In the next subsections, we choose a particular configuration for the strain and the external magnetic field and explore the differences between these three interactions.
@@ III.1 Strain tensor
with $\lambda$ and $\mu$ are the Lamé coefficients and $\theta=u^{i}_{i}$ is the trace of the strain tensor. From the ansatz employed in Eq.(19), we obtain the strain tensor as
@@
The form of the strain tensor in Eq.(23) shows that the deformations undergone by the surface are isotropic and concentrated aroun the catenoid throat. The position–dependent Fermi velocity tensor $v_{ij}$ can be written as
@@
$v(u)=v_{0}\left[1-\frac{\beta}{2}\frac{\sigma(u)}{\lambda+\mu}\right].$ (25)
@@
$\Gamma_{u}$ $=$ $-\frac{\beta}{4}\frac{\sigma^{\prime}(u)}{\lambda+\mu},$ (26)
@@ III.2 Magnetic vector potential
Now, let us see how the external magnetic field $\vec{B}$ modifies the Hamiltonian. For an uniform magnetic field $\vec{B}=B\hat{k}$, the vector potential $\vec{A}$ is given by $\vec{A}=\frac{1}{2}\vec{B}\times\vec{r}$. Using the coordinates system in Eq.(1), we obtain
@@
For $R=0$, the expression for the vector potential in Eq.(29) reduces to $\vec{A}=\pm\frac{B}{2}u\hat{e}_{2}$, as found in Ref.(diracmagnetic1 ). For $u\gg R$, it yields to $A^{2}\approx\frac{B}{2}u$, as in a conical surface diracmagnetic1 and on the flat plane diracplanar . In addition, at $u=0$, it has a finite value $A^{2}=\frac{BR}{2}$. Since $\vec{e}_{\phi}=\sqrt{R^{2}+u^{2}}\hat{e}_{2}$, the component of $\vec{A}$ in coordinates is given by $A^{\phi}=\frac{B}{2}$. Accordingly, the contravariant component $A_{\phi}=g_{\phi\phi}A^{\phi}$, we have
@@ IV Effective Hamiltonian
where $\bar{\beta}=\frac{\beta}{4\lambda+\mu}$. Notice that the strain vector and the spinor connection modifies the Dirac equation leading to a canonically momentum on the $u$ of form
@@
The expression in Eq.(31) depends only on the coordinate $u$. The symmetry of the Hamiltonian with respect to the angular $\phi$ variable is the result of the surface axial symmetry. Thus, the wave function should also inherit this symmetry. In fact, consider the angular momentum operator with respect to the $\mathrm{z}$ axis, $\hat{L}_{z}=-i\hbar\frac{\partial}{\partial\phi}$ such that, $\hat{L}_{z}\psi=\hbar l\psi$, where $l$ is the orbital angular momentum with respect to the $\mathrm{z}$ axis. For a non–relativistic and spinless electron on the graphene wormhole, an axissymmetric wave function can be written as $\psi(u,\phi)=e^{il\phi}\psi(u)$ euclides . However, as it is well–known for the relativistic electron, $\hat{L}_{z}$ no longer commute with $\mathcal{H}_{D}$, although the total angular momentum operator along the $z$ direction $\hat{J}_{z}=\hat{L}_{z}+\hat{S}_{z}$ does diracplanar . Since $\hat{S}_{3}=\frac{\hbar}{4}[\gamma^{1},\gamma^{2}]$, then the spin operator with respect to the $\mathrm{z}$ axis is given by
@@
Since the effective Hamiltonian in Eq.(38) has parity violating terms steaming from the geometric connection and the strain vector, we can not assume that the spinor $\psi=\left(\begin{array}[]{cc}\psi_{1}\\ \psi_{2},\end{array}\right)$ is parity invariant. This parity violation of the Dirac spinor on the graphene wormhole is in contrast with the relativistic electron on a flat plane diracplanar ; diracplanar2 .
In Eq.(38), the position–dependent velocity function $v(u)$ multiplies the partial derivative $\partial_{u}$. By writing $v\frac{d}{du}=\frac{d}{d\zeta}$, for $v(u)$ given by Eq.(25), we have $\zeta=u-\bar{\beta}R\,\tanh^{-1}\left(\frac{\sqrt{u}u}{\sqrt{\bar{\beta}-2}R}\right)$. Unfortunately, this relation analytically can not be inverted, seeking to rewrite Eq. (38) in terms of variable $\zeta$. Nevertheless, a graphic analysis reveals only a small difference between $\zeta$ and $u$ as shown in Fig. 12. Analogously, for the sake of simplicity, we adopt $\zeta\approx u$ from now on.
@@
The effective Hamiltonian in Eq.(41) shows clearly the distinctive interaction terms steaming from the strain $\bar{\beta}\sigma^{\prime}$, from the geometric connection $\frac{u}{2(R^{2}+u^{2})}$, centrifugal term $\frac{m}{\sqrt{R^{2}+u^{2}}}$ and the electromagnetic coupling $\frac{eA_{\phi}}{\sqrt{R^{2}+u^{2}}}$. In the next section, we explore the effects of each interaction.

### v2

@@ Abstract
In this work, we explore the strain and curvature effects on the electronic properties of a curved graphene structure, called the graphene wormhole. The electron dynamics is described by a massless Dirac fermion containing position–dependent Fermi velocity. In addition, the strain produces a pseudo–magnetic vector potential to the geometric coupling. For an isotropic strain tensor, the decoupled components of the spinor field exhibit a supersymmetric (SUSY) potential, depending on the centrifugal term and the external magnetic field only. In the absence of an external magnetic field, the strain yields an exponentially damped amplitude, whereas the curvature leads to a power–law damping of the wave function. The spin–curvature coupling breaks the chiral symmetry between the upper and the lower spinor component, which leads to the increasing of the wave function on either upper or lower region of the wormhole, i.e., depending on the spin number. By adding a uniform magnetic field, the effective potential exhibits an asymptotic quadratic profile and a spin–curvature barrier near the throat. As a result, the bound states (Landau levels) are confined around the wormhole throat showing an asymmetric and spin–dependent profile.
@@ I Introduction
Two dimensional materials, such as graphene geim , silicene silicene and phosphorene phosphorene , have been the subject of intense investigation due to their outstanding properties. Beyond the remarkable mechanical katsnelson and electronic properties Novoselov2004 ; electronic , graphene can also be seen as a table–top laboratory for relativistic physics. Indeed, since the conduction electrons are effectively described as massless Dirac fermions, relativistic effects such as zitterbewegun zitter , Klein tunneling klein and atomic collapse collapse have been observed. Since the graphene layer can assume a curved shape, the curvature effects might lead to new interesting relativistic effects, such as the Hawking–Unruh effect hawking ; hawking2 .
@@
The study of a Dirac fermion confined into a two dimensional surface was initially addressed in Ref.BJ and further developments were provided afterwards diracsurface ; diracsurface2 ; diracsphere . For a relativistic fermion intrinsically living on a curved surface, a physical realization was found for conducting electrons on two dimensional carbon–based structures, as the fullerenes diracintrinsic ; diracintrinsic1 , carbon nanotubes saito and graphitic cones cone . In graphene, the massless Dirac equation in curved spaces was studied in a variety of shapes, such as the localized gaussian bump contijo , the cone furtado , a helical graphene ribbon atanasov ; watanabe , a corrugated plane corrugated , a Möbius ring mobius1 ; mobius2 , a torus ozlem among others. The surface curvature produces a spin–curvature coupling which leads to a geometric Aharonov–Bohm–like effect geometricphase , a modified spin–orbit coupling geometricsoc ; geometricsoc2 and a geometric spin–Hall effect geometricmonopole .
@@
In addition to the curvature, the deformations of the graphene layer modify the effective Dirac fermion dynamics as well, producing the so–called pseudo–magnetic fields ribbons . This vector potential stems from the strain tensor defined by the deformations of the graphene layer and the pseudo–magnetic term comes from the coupling to the Dirac fermion; it is similar to the minimal coupling to a magnetic field gaugestrain . The strain applied to graphene can mimic a strong magnetic field strainstrong and leads to important applications strainappli . From the strain tensor, an effective Hamiltonian for the Dirac fermion was derived using the tight–binding approach vozmediano , containing an anisotropic and position–dependent Fermi velocity and a pseudo–magnetic strain vector in the continuum limit. Since the strain may contain both in–plane and out–of–plane components, the effective Dirac Hamiltonian was extended in order to encompass all the stretching and bending effects vozmediano2 . In addition, the quantum field interaction of the effective Dirac fermion and the strain was discussed in sinner ; gaugegrapheneqft .
@@
An interesting curved graphene structure is the so–called graphene wormhole, where two flat graphene layers are connected by a carbon nanotube wormhole . Since its shape (cylinder) has a non–vanishing mean curvature, the discontinuity of the curvature at the graphene–nanotube junction leads to modifications of the energy spectrum and the possibility of localized states close to it graphenejunction . Although Dirac fermions on the upper and lower layers are free states (non–normalizable), the curvature of the nanotube allows the existence of normalizable zero–modes confined at the radius of the wormhole picak ; wormhole3 . In order to avoid the discontinuity at the junction, a smooth graphene wormhole was proposed considering a continuum and asymptotic flat catenoid surface dandoloff ; euclides . The negative curvature of the catenoid leads to a repulsive spin–curvature coupling near the wormhole throat, allowing only the zero–mode as a localized state around the throat wormhole4 ; ozlem2 ; wormhole5 .
@@
In this work, we consider the effects of the curvature and the strain on the effective Dirac fermion living in a catenoid–shaped graphene wormhole. We extend the effective Hamiltonian obtained in vozmediano to a curved surface, introducing the usual spin–connection coupling. We explore the different effects driven by the curvature, isotropic strain and an external magnetic field. Since the surface is asymptotically flat, the lattice deformation which produces the curvature and strain should be concentrated around the throat. The strain leads to a vector potential along the surface meridian, whereas the spin–curvature coupling points in the angular direction. Moreover, the strain vector potential provides an exponential damping of the wave function, whereas the curvature leads to a power–law decay. By adopting the so-called supersymmetric quantum mechanic-like approach ozlem2 , the spinor components exhibit a chiral symmetry breaking. Indeed, the upper component has its probability density enhanced near the wormhole throat in the upper layer, whereas the lower component is enhanced in the lower layer. The ground state zero mode also exhibits this chiral behaviour, since it is exponentially damped either in the upper or in the lower layer depending on the total angular momentum. By applying a uniform magnetic field, the Landau levels are also modified by the curved geometry and strain, leading to asymmetric localized states near the throat.
@@
This work is organized as follows. In section (II), we provide a brief review of the geometry of the catenoid-shaped graphene wormhole. In section (III) we present the effective Hamiltonian containing the strain, curvature and external magnetic field interactions. The section (IV) is devoted to the symmetries of the effective Hamiltonian and in the section (V) we employ the SUSY-QM approach in order to investigate the effects of each interaction. Finally, additional discussion and perspectives are given in the section (VI).
@@ II Graphene wormhole geometry
In this section, we define the surface of the graphene wormhole and describe some of its most important properties. We consider a smooth surface connecting an upper to the lower layer (flat planes). For this purpose, we choose a catenoid shaped surface. In other words, the catenoid surface can be described in coordinates by euclides ; ozlem2
@@
Remember that they modify the Fermi velocity by turning it into a position dependent configuration. Moreover, from the dreinbeins, we can define the moving frame $\theta^{a}=e^{a}_{\mu}\mathrm{d}x^{\mu}$, where, in the catenoid, takes the form $\theta^{0}=\mathrm{d}t$, $\theta^{1}=\mathrm{d}u$, $\theta^{2}=\sqrt{R^{2}+u^{2}}\mathrm{d}\phi$. From the torsion–free condition, i.e., $T^{a}=\mathrm{d}\theta^{a}+\omega^{a}_{b}\wedge\theta^{b}=0$, the only non–vanishing one–form connection coefficient $\omega^{a}_{b}=\Gamma_{cb}^{a}\theta^{c}$ is given by
@@
The curvature 2–form $R^{a}_{b}=d\omega^{a}_{b}+\omega^{a}_{c}\wedge\omega^{c}_{b}$ has only one non–vanishing component, namely $R^{2}_{1}=-\frac{R^{2}}{(R^{2}+u^{2})^{2}}\theta^{2}\wedge\theta^{1}$. Accordingly, the gaussian curvature $K=\delta^{ab}R_{ab}$ has the form
@@ III Strain Hamiltonian
After a brief review of the main geometric properties of the graphene wormhole, let us now discuss the effective Hamiltonian containing the strain and curvature effects on the electron. We follow closely the Ref.(vozmediano ) where the most general effective Hamiltonian was derived. The effective Hamiltonian for a Dirac fermion constrained to a flat surface under the influence of the strain and an external magnetic field was found in Ref. (vozmediano ). We propose a generalization of the effective Hamiltonian of Ref.(vozmediano ) in the continuum limit for curved surfaces in the form
@@
The definition of the curved sigma matrices employed in eq.(12) ensures that these matrices do not depend on the particular choice of coordinates of the surface (surface covariance). It is worthwhile to mention that the definitions of the velocity tensor in eq.(24) and the curved Pauli matrices in eq.(12) lead to a position and direction dependent Dirac kinetic term $H_{1}=v_{i}^{j}\sigma^{i}\partial_{j}$.
@@
The curved $\gamma^{\mu}$ matrices are related to the flat $\gamma^{a}$ ones by the dreinbeins $e^{a}_{\mu}$, i.e., $\gamma^{\mu}=e^{\mu}_{a}\gamma^{a}$. The dreinbeins are defined as $g_{\mu\nu}=e_{\mu}^{a}e_{\nu}^{b}$. In $(2+1)$ dimension, we can adopt the following representation for the flat Dirac $\gamma^{a}$ matrices $\gamma_{0}=\sigma_{3}$, $\gamma_{1}=-i\sigma_{2}$ and $\gamma_{2}=-i\sigma_{1}$ mobius2 ; ozlem ; ozlem2 . Thus, the curved Dirac matrices on the wormhole graphene surface have the form
@@
From the connection 1–form in eq.(7), only $\omega^{2}_{1}$ is non-zero. Thus, the only non–vanishing component of the spinor connection $\Omega_{\mu}$ is
@@
Note that, since $-\infty<u<\infty$, the geometric spinor potential in eq.(18) is an odd function under parity. For $R=0$ or for $u\rightarrow\pm\infty$, the spinor connection becomes constant $\Omega_{\phi}=\frac{i}{2}\sigma_{3}$, as obtained in the graphitic cone furtado . the geometric connection is constant, as found for conic surfaces furtado . It is worth mentioning that, due to the resemblance of the spinor and gauge field couplings, the spinor potential is sometimes interpreted as a kind of a pseudo–magnetic potential stemming from the curved geometry contijo .
Furthermore, the strain produces two different potentials on the Dirac electron. The first potential, due to the strain in eq.(11), is a vector potential, while the second one in eq.(18) is a spinorial potential depending on the surface connection. In the next subsections, we choose a particular configuration for the strain and the external magnetic field and explore the differences between these three interactions.
@@ III.1 Strain tensor
where $\lambda$ and $\mu$ are the Lamé coefficients and $\theta=u^{i}_{i}$ is the trace of the strain tensor. From the ansatz employed in eq.(19), we obtain the strain tensor as
@@
The form of the strain tensor in eq.(23) shows that the deformations undergone by the surface are isotropic and concentrated around the catenoid throat. The position–dependent Fermi velocity tensor $v_{ij}$ can be written as
@@
$v(u)=v_{0}\left[1-\frac{\beta}{2}\frac{\sigma(u)}{(\lambda+\mu)}\right].$ (25)
@@
$\Gamma_{u}$ $=$ $-\frac{\beta}{4}\frac{\sigma^{\prime}(u)}{(\lambda+\mu)},$ (26)
@@ III.2 Magnetic vector potential
Now, let us see how the external magnetic field $\vec{B}$ modifies the Hamiltonian. For a uniform magnetic field $\vec{B}=B\hat{k}$, the vector potential $\vec{A}$ is given by $\vec{A}=\frac{1}{2}\vec{B}\times\vec{r}$. Using the coordinates system in eq.(1), we obtain
@@
For $R=0$, the expression for the vector potential in eq.(29) reduces to $\vec{A}=\pm\frac{B}{2}u\hat{e}_{2}$, as found in Ref.(diracmagnetic1 ). For $u\gg R$, it yields $A^{2}\approx\frac{B}{2}u$, as in a conical surface diracmagnetic1 and on the flat plane diracplanar . In addition, at $u=0$, it has a finite value $A^{2}=\frac{BR}{2}$. Since $\vec{e}_{\phi}=\sqrt{R^{2}+u^{2}}\hat{e}_{2}$, the component of $\vec{A}$ in coordinates is given by $A^{\phi}=\frac{B}{2}$. Accordingly, for the covariant component $A_{\phi}=g_{\phi\phi}A^{\phi}$, we have
@@ IV Effective Hamiltonian
where $\bar{\beta}=\frac{\beta}{4(\lambda+\mu)}$. Notice that the strain vector and the spinor connection modifies the Dirac equation leading to a canonical momentum along the $u$ direction of form
@@
The expression in eq.(31) depends only on the coordinate $u$. The symmetry of the Hamiltonian with respect to the angular $\phi$ variable is the result of the surface axial symmetry. Thus, the wave function should also inherit this symmetry. In fact, consider the angular momentum operator with respect to the $\mathrm{z}$ axis, $\hat{L}_{z}=-i\hbar\frac{\partial}{\partial\phi}$ such that, $\hat{L}_{z}\psi=\hbar l\psi$, where $l$ is the orbital angular momentum with respect to the $\mathrm{z}$ axis. For a non–relativistic and spinless electron on the graphene wormhole, an axisymmetric wave function can be written as $\psi(u,\phi)=e^{il\phi}\psi(u)$ euclides . However, as it is well–known for the relativistic electron, $\hat{L}_{z}$ no longer commute with $\mathcal{H}_{D}$, although the total angular momentum operator along the $z$ direction $\hat{J}_{z}=\hat{L}_{z}+\hat{S}_{z}$ does diracplanar . Since $\hat{S}_{3}=\frac{\hbar}{4}[\gamma^{1},\gamma^{2}]$, then the spin operator with respect to the $\mathrm{z}$ axis is given by
@@
The effective Hamiltonian in eq.(38) has parity-even and parity-odd potentials.
In eq.(38), the position–dependent velocity function $v(u)$ multiplies the partial derivative $\partial_{u}$. By writing $v\frac{d}{du}=\frac{d}{d\zeta}$, for $v(u)$ given by eq.(25), we have $\zeta=u-\frac{\bar{\beta}R}{\sqrt{2(\bar{\beta}-2)}}\,\tanh^{-1}\left(\frac{\sqrt{2}u}{R\sqrt{(\bar{\beta}-2)}}\right)$. Unfortunately, this relation can not be inverted analytically, as one seeks to rewrite eq. (38) in terms of variable $\zeta$. However, a graphic analysis reveals that, for $0\leq\bar{\beta}\leq 1$, there is only a small difference between $\zeta$ and $u$, as shown in fig. (12) for $\bar{\beta}=1$. Indeed, taking the limit $\bar{\beta}\rightarrow 1$, the function $\zeta$ becomes $\zeta=u+\frac{R}{\sqrt{2}}\tan^{-1}{\sqrt{2}u}$. Note that for $u\rightarrow\pm\infty$, $\zeta-u\rightarrow\pm\frac{R}{\sqrt{2}}$, and thus $\frac{\zeta-u}{u}\rightarrow 0$. Therefore, near the origin $\zeta\approx u$ and asymptotically $\zeta$ is only $u$ shifted by a constant amount, as we can also see by the graphic in fig.(12). For $21<\bar{\beta}<$ 2, the difference $\zeta-u$ increases, though it is still finite and $\frac{(\zeta-u)}{u}\rightarrow 0$, as $u\rightarrow\infty$. At $\bar{\beta}=2$ the function $\zeta$ is not defined and for $\bar{\beta}>2$ the function $\zeta$ is only defined in the interval $|u|<\frac{R}{\sqrt{2}}\sqrt{\bar{\beta}-2}$. Since we are interested in solutions for all $-\infty<u<\infty$ and we are considering small strain effects, i.e., $\bar{\beta}\sigma_{0}\leq 1/2$, then we restrict ourselves to the interval $0\leq\bar{\beta}\leq 1$. Accordingly, since the position-dependence of the Fermi velocity is small for $0\leq\bar{\beta}\leq 1$, for the sake of simplicity we adopt $\zeta\approx u$ from now on.
@@
The effective Hamiltonian in eq.(41) shows clearly the distinctive interaction terms arising from the strain $\bar{\beta}\sigma^{\prime}$, from the geometric connection $\frac{u}{2(R^{2}+u^{2})}$, centrifugal term $\frac{m}{\sqrt{R^{2}+u^{2}}}$ and the electromagnetic coupling $\frac{eA_{\phi}}{\sqrt{R^{2}+u^{2}}}$. In the next section, we explore the effects of each interaction.